OpenStudy (anonymous):

what percent of 150 is 24.6

OpenStudy (anonymous):

Divide 24.6 by 150 and then multiply by 100

OpenStudy (anonymous):

That's finding a percentage.

OpenStudy (anonymous):

Do you get it?

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

\[24.6 = \frac{x }{ 100 }*150\\\\ 24.6= \frac{ x }{ 2 }*3\\\\ x\frac{ 2*24.6 }{3 }\\\\ x= 16.4 \% \]